OpenStudy (anonymous):

write 5/9 with a denominator of 45

OpenStudy (anonymous):

well if you want to write itthe denominator of 45 you first need to write it down like this 5/9 = 1/45 5 * 5 = 25 9 * 5 = 45 so the answer would be 25/45